Tohban Report 2015-04-08
From RHESSI Wiki
|Start Date:||01 April 2015|
|End Date:||08 April 2015|
|List all reports|
While solar activity has been mostly low this week, there have been a few isolated large C and low M flares. (The single M flare doesn't seem to be accounted for in the table below.)
How many GOES flares occurred?
Flares above B, C, M, X class were 11 11 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 8 10 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 140 / 22 over the time range 01-Apr-15 08-Apr-15
The SSR is still not emptying at the end of the pass set. This could be due to residual ground station issues (i.e. missing some Wallops passes, and a recent glitch at BGS). However, it also might be due to excess data from a noisy detector. We will watch to see how the recent threshold changes affect this issue. The SSR is not at a worrying level (ending the pass set at ~10% or less).
A spinup was performed on April 1. RHESSI was spun up from 14.6 to 15.0 RPM between 2015-18:52 and 20:22 UTC.
There was one data gap:
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2015-04-06T20:10:00.000 -- 2015-04-06T23:30:00.000 12000.000
This is being investigated to see if we have the data on the ground and if it needs reprocessing.
Cryocooler power: It was determined that the cryocooler experiment performed the previous week (i.e. turning up the power to 75W) had not caused a significant enough change in the cold tip temperature (only 0.6 degree total) and was causing some potential for damaging problems, including damaging the cryocooler (though note we don't think any damage was caused by this test), running too low on the battery voltage (due to the increased power usage), and additional noise on the spacecraft bus that may have caused an ATS load not to take. It was noted that what we are seeing from the cryocooler, i.e. decreased return (cooling) on the investment (power) is indicative of approaching the regime where we're almost out of room for the cryocooler piston to move without hitting the stop. In light of these risk factors, and the fact that we didn't gain much in terms of temperatures, it was decided to return the cryocooler power to 71W.
That change was made April 1 (no joke!) at 21:59 UTC.
Detector 9's rear fast rates have increased to the point where they are eating up all livetime, and as a result the rear segment of detector 9 is returning no actual events. The fast threshold is already as high as it can go, so there is no obvious way to ameliorate this problem. For the time being (and the foreseeable future) the rear segment of D9 is not usable.
Very low livetime was observed in D8 rear and D6 & D9 fronts, so fast thresholds were changed on all of these on April 7. While some improvements were seen, high slow rates were also observed, so changes to the slow thresholds might instead be needed. (Note that this will cut into the energy range of recorded events for the front segments.) The next tohban will asses the need for further changes.
15-097-19:56:35 /IDPUTABLE6 OFFSET=FRONTFASTDAC 15-097-19:56:45 /IDPLOAD VALUE=0x80 ; was 0x70
15-097-19:57:24 /IDPUTABLE8 OFFSET=REARFASTDAC 15-097-19:57:37 /IDPLOAD VALUE=0x70 ; was 0x65
15-097-19:58:06 /IDPUTABLE9 OFFSET=FRONTFASTDAC 15-097-19:58:14 /IDPLOAD VALUE=0x30 ; was 0x22
The behavior of these detectors before/after the changes:
- D6 front: high fast rates / extremely low livetime before the change. The threshold change improved, but not much, so a higher raise may be necessary.
- D9 front: livetime at ~90% before and after the change; no noticeable difference.
- D8 rear: Extremely low livetime (almost zero) before the change. Unclear whether raising the threshold helped. Further changes may be necessary.
|Night time data (fronts)||plus/minus 4 min|
|Night time data (rears)||plus/minus 4 min|
|Require extra passes?||No, but maybe adding Weilheim next week.|
|Requirement for moving pointer?||No|
|Attenuator operation||Normal (actually, no motions, but this seems ok!)|